find an equation of the line tangent to the curve y=3x2-1 and parallel to the line 2x-y+3


1
Expert's answer
2021-03-31T15:57:52-0400

Let us find an equation of the line tangent to the curve "y=3x^2-1" and parallel to the line "2x-y+3=0." The equation of the tangent line at point "x_0" is "y=y(x_0)+y'(x_0)(x-x_0)". Since the tangent line and the line "y=2x+3" are parallel, their slopes coinside, that is "y'(x_0)=2". Taking into account that "y'(x)=6x", we conclude that "6x_0=2", and hence "x_0=\frac{1}{3}", "y(\frac{1}{3})=3\frac{1}{9}-1=-\frac{2}{3}." It follows that the equation of the tangent line is the following:


"y=-\frac{2}{3}+2(x-\frac{1}{3})" or "y=2x-\frac{4}{3}."